Evaluate the integral of cos2x dx/(2+5 sin 2x)⁴


1
Expert's answer
2021-03-26T04:30:55-0400
"\int\dfrac{\cos(2x)}{(2+5\sin(2x))^4}dx"

Use "u" -substitution


"u=2+5\sin(2x), du=10\cos(2x)dx"

"\int\dfrac{\cos(2x)}{(2+5\sin(2x))^4}dx=\dfrac{1}{10}\int\dfrac{du}{u^4}dx"

"=-\dfrac{1}{30u^3}+C=-\dfrac{1}{30(2+5\sin(2x))^3}+C"

"\int\dfrac{\cos(2x)}{(2+5\sin(2x))^4}dx=-\dfrac{1}{30(2+5\sin(2x))^3}+C"